In 2016, Western Maryland Company paid $3,000,000 for a mining tract with reserves of 100,000 tons of ore, which the company planned to mine over a 20-year period. The company spent $500,000 developing the mine. Western plans to spend $200,000 for reclamation when mining has been completed, after which the land will have an estimated value of $600,000. In 2016, Western mined and sold 8,000 tons of ore. What depletion should Western record for the year?

A) $155,000

B) $296,000

C) $248,000

D) $280,000
Answer #1

Depletion rate per ton

= (Cost - Salvage value) /no of total tons

= (3,000,000 + 500,000 + 200,000 - 600,000)/100,000

= 3,100,000/100,000

= 31 pet ton

Depletion expense for the year = 31 * 8,000

= 248,000